I need help with this question please and thank you

Problem3. Solve the initial value problem y = 5y + e -*y2 y(O)= 2
Answer #1

soli - Given y= 5y + e 22 j2 - 0 460]=2 yl-5y = 21,527 Equation ☺ is Bernoullis Equation that is dy + py = Q yh I de Q = ē nO da Equation ® is linear differential Equation. that is du + pv=a. where = -15 a = 3 eat Integrating factor (I.F) = ( pda =ľ in Egn ® substitute y = [-3ē + 139 ent] 17 yo = 11-3222 +1390152